# Complexity and curvature of pairs of Burch modules and ideals

### v1

@@ 3. Main results
(1) By Lemma 2.16, there exists an $M$-regular element $x\in\mathfrak{m}$ such that $k$ is a direct summand of $M/xM$. Therefore $\operatorname{cx}_{R}(N)=\operatorname{tcx}_{R}(k,N)\leqslant\operatorname{tcx}_{R}(M/xM,N)\leqslant\operatorname{tcx}_{R}(M,N)$, where the last two inequalities are obtained from Lemma 2.10.(1) and [DGS24, Lem. 3.12.(2)] respectively. Similarly, one obtains $\operatorname{curv}_{R}(N)\leqslant\operatorname{tcurv}_{R}(M,N)$.
@@
(1) By Lemma 2.16, there exists an $M$-regular element $x$ of $R$ such that $k$ is a direct summand of $M/xM$. Therefore $\operatorname{inj\,cx}_{R}(N)=\operatorname{cx}_{R}(k,N)\leqslant\operatorname{cx}_{R}(M/xM,N)\leqslant\operatorname{cx}_{R}(M,N)$, where the last two inequalities are obtained from Lemma 2.10.(3) and [DGS24, Lem. 3.13.(2)] respectively. Similarly, one can obtain $\operatorname{inj\,curv}_{R}(N)\leqslant\operatorname{curv}_{R}(M,N)$.
@@
(1) By Lemma 2.16, there exists an $N$-regular element $x$ of $R$ such that $k$ is a direct summand of $N/xN$. Therefore $\operatorname{cx}_{R}(M)=\operatorname{cx}_{R}(M,k)\leqslant\operatorname{cx}_{R}(M,N/xN)\leqslant\operatorname{cx}_{R}(M,N)$, where the last two inequalities are obtained by using Lemma 2.10.(5) and [DGS24, Lem. 3.13(1)] respectively. The inequality for curvature can be obtained using similar arguments.
@@
All hypotheses and conclusions are preserved under passing to completion (for the Burch property, see Lemma 2.15). Hence, we may assume that $R$ is complete. Let $N$ be a Burch submodule of an $R$-module $X$. Since $R$ is complete and CM, there exist a finitely generated $R$-module $L$ of finite injective dimension and an injective homomorphism $h:X\to L$ (see [lw, 11.8 and 11.17]). By Lemma 2.19, $h(N)$ is a Burch submodule of $L$. Note that $\operatorname{Ext}^{\gg 0}_{R}(M,L)=0$ (i.e., $\operatorname{cx}_{R}(M,L)=0$), and $h(N)\cong N$. Therefore, by Proposition 3.9, $\operatorname{cx}_{R}(M)\leqslant\operatorname{cx}_{R}(M,h(N))=\operatorname{cx}_{R}(M,N)$, and similarly $\operatorname{curv}_{R}(M)\leqslant\operatorname{curv}_{R}(M,N)$. ∎
@@
Set $X:=\operatorname{\Omega}^{d}_{R}(M)$. Then $X$ is MCM, and $\operatorname{Ext}^{n+d}_{R}(M,N)\cong\operatorname{Ext}_{R}^{n}(X,N)$ for all $n\geqslant 1$. Consider an MCM approximation (cf. [lw, 11.8 and 11.17]) of $N$, i.e., a short exact sequence $0\to L\to Y\to N\to 0$, where $L$ has finite injective dimension, and $Y$ is MCM. Hence, the desired isomorphisms follow because $\operatorname{Ext}_{R}^{n}(X,L)=0$ for all $n\geqslant 1$. Note that $Y$ locally has finite injective dimension on punctured spectrum as both $L$ and $N$ have the same property. ∎
@@
Set $d:=\dim(R)$. As in the proof of Theorem 3.10, passing to completion, we may assume that $R$ is complete, and $N$ locally has finite injective dimension on the punctured spectrum. Hence, since $R$ is CM, it admits a canonical module, say $\omega$. Set $(-)^{\dagger}:=\operatorname{Hom}_{R}(-,\omega)$. By Lemma 3.11, there exist MCM $R$-modules $X$ and $Y$ such that $Y$ locally has finite injective dimension on the punctured spectrum, and $\operatorname{Ext}^{n+d}_{R}(M,N)\cong\operatorname{Ext}_{R}^{n}(X,Y)\cong\operatorname{Ext}_{R}^{n}(Y^{\dagger},X^{\dagger})$ for all $n\geqslant 1$, where the last isomorphism is by duality of dualizing module, see, e.g., [cluster, 3.14 and 3.15]. Since $Y$ is MCM and locally has finite injective dimension on the punctured spectrum, for every non-maximal prime ideal $\mathfrak{p}$ of $R$, the localization $Y_{\mathfrak{p}}$ is a finite direct sum of copies of $\omega_{\mathfrak{p}}$. Thus, $Y^{\dagger}$ is locally free on the punctured spectrum of $R$. Therefore, by [DGS24, Lem. 3.6], there exists $h>0$ such that $0=\mathfrak{m}^{h}\operatorname{Ext}_{R}^{n}(Y^{\dagger},X^{\dagger})=\mathfrak{m}^{h}\operatorname{Ext}_{R}^{n+d}(M,N)$ for all $n\gg 0$. Hence, the proposition follows from Theorem 3.10. ∎
@@
In general, it may happen that $R$ is CM with isolated singularity, but $\widehat{R}$ is locally not even Gorenstein on the punctured spectrum, see [heit, Main Theorem]. In the following remark, we note a class of local rings whose completions are locally Gorenstein on the punctured spectrum.
@@
    Suppose that $R$ is CM of dimension $1$. Then it is well known that $\widehat{R}$ is locally Gorenstein on the punctured spectrum if and only if $R$ admits a canonical ideal. Indeed, in view of [agl, Prop. 2.7], $R$ admits a canonical ideal if and only if the total ring of fractions Q(R^)Q($\widehat${R}) is Gorenstein if and only if $\widehat{R}$ is locally Gorenstein at all minimal primes of $\widehat{R}$, where the last equivalence follows from [stacks-project, Lem. 10.25.4] and the fact that the direct product of finitely many rings is Gorenstein if and only if all the components are Gorenstein.
@@
First, assume that $\operatorname{G-dim}_{R}(X)<\infty$. Then, in view of 2.26, there exists an $R$-module $L$ of finite projective dimension and an injective homomorphism $h:X\to L$. By Lemma 2.19, $h(M)$ is a Burch submodule of $L$. Note that $\operatorname{Tor}^{R}_{\gg 0}(L,N)=0=\operatorname{Ext}^{\gg 0}_{R}(L,N)$, and $h(M)\cong M$. Therefore, we are done by Propositions 3.7 and 3.8.
@@
    The hypotheses $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(M,N)=0$ and/or $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(L,N)=0$ for some $h\geqslant 1$ in Propositions 3.1, 3.7 and Theorems 3.2, 3.15.(1) hold if $N$ locally has finite projective dimension on the punctured spectrum; see [DGS24, Lem. 3.6.(2)].
@@
    The hypotheses $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,N)=0$ and/or $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,L)=0$ for some $h\geqslant 1$ in Propositions 3.3, 3.5, 3.9 and Theorems 3.6, 3.10, 3.15.(2) hold if $M$ locally has finite projective dimension on the punctured spectrum; see [DGS24, Lem. 3.6.(1)].
@@
Although different in spirit from the rest of the results in this article, the following proposition demonstrates the usefulness of Lemma 2.19 by improving an injective dimension test result from [GS24, Cor. 5.4].
@@
(1) In view of 2.26, there exist an $R$-module $L$ of finite projective dimension and an injective homomorphism $h:X\to L$. By Lemma 2.19, $h(M)$ is a Burch submodule of $L$. Since $n>\operatorname{depth}(R)\geqslant\operatorname{pd}_{R}(L)$, it follows that $\operatorname{Ext}_{R}^{i}(L,N)=0$ for $i=n,n+1$. Since $h(M)\cong M$, by [GS24, Thm. 5.3], $\operatorname{id}_{R}(N)<\infty$.

### v2

@@ 3. Main results
(1) By Lemma 2.16, there exists an $M$-regular element $x\in\mathfrak{m}$ such that $k$ is a direct summand of $M/xM$. Therefore $\operatorname{cx}_{R}(N)=\operatorname{tcx}_{R}(k,N)\leqslant\operatorname{tcx}_{R}(M/xM,N)\leqslant\operatorname{tcx}_{R}(M,N)$, where the last two inequalities are obtained from Lemma 2.10.(1) and [12, Lem. 3.12.(2)] respectively. Similarly, one obtains $\operatorname{curv}_{R}(N)\leqslant\operatorname{tcurv}_{R}(M,N)$.
@@
(1) By Lemma 2.16, there exists an $M$-regular element $x$ of $R$ such that $k$ is a direct summand of $M/xM$. Therefore $\operatorname{inj\,cx}_{R}(N)=\operatorname{cx}_{R}(k,N)\leqslant\operatorname{cx}_{R}(M/xM,N)\leqslant\operatorname{cx}_{R}(M,N)$, where the last two inequalities are obtained from Lemma 2.10.(3) and [12, Lem. 3.13.(2)] respectively. Similarly, one can obtain $\operatorname{inj\,curv}_{R}(N)\leqslant\operatorname{curv}_{R}(M,N)$.
@@
(1) By Lemma 2.16, there exists an $N$-regular element $x$ of $R$ such that $k$ is a direct summand of $N/xN$. Therefore $\operatorname{cx}_{R}(M)=\operatorname{cx}_{R}(M,k)\leqslant\operatorname{cx}_{R}(M,N/xN)\leqslant\operatorname{cx}_{R}(M,N)$, where the last two inequalities are obtained by using Lemma 2.10.(5) and [12, Lem. 3.13(1)] respectively. The inequality for curvature can be obtained using similar arguments.
@@
All hypotheses and conclusions are preserved under passing to completion (for the Burch property, see Lemma 2.15). Hence, we may assume that $R$ is complete. Let $N$ be a Burch submodule of an $R$-module $X$. Since $R$ is complete and CM, there exist a finitely generated $R$-module $L$ of finite injective dimension and an injective homomorphism $h:X\to L$ (see [17, 11.8 and 11.17]). By Lemma 2.19, $h(N)$ is a Burch submodule of $L$. Note that $\operatorname{Ext}^{\gg 0}_{R}(M,L)=0$ (i.e., $\operatorname{cx}_{R}(M,L)=0$), and $h(N)\cong N$. Therefore, by Proposition 3.9, $\operatorname{cx}_{R}(M)\leqslant\operatorname{cx}_{R}(M,h(N))=\operatorname{cx}_{R}(M,N)$, and similarly $\operatorname{curv}_{R}(M)\leqslant\operatorname{curv}_{R}(M,N)$. ∎
@@
Set $X:=\operatorname{\Omega}^{d}_{R}(M)$. Then $X$ is MCM, and $\operatorname{Ext}^{n+d}_{R}(M,N)\cong\operatorname{Ext}_{R}^{n}(X,N)$ for all $n\geqslant 1$. Consider an MCM approximation (cf. [17, 11.8 and 11.17]) of $N$, i.e., a short exact sequence $0\to L\to Y\to N\to 0$, where $L$ has finite injective dimension, and $Y$ is MCM. Hence, the desired isomorphisms follow because $\operatorname{Ext}_{R}^{n}(X,L)=0$ for all $n\geqslant 1$. Note that $Y$ locally has finite injective dimension on punctured spectrum as both $L$ and $N$ have the same property. ∎
@@
Set $d:=\dim(R)$. As in the proof of Theorem 3.10, passing to completion, we may assume that $R$ is complete, and $N$ locally has finite injective dimension on the punctured spectrum. Hence, since $R$ is CM, it admits a canonical module, say $\omega$. Set $(-)^{\dagger}:=\operatorname{Hom}_{R}(-,\omega)$. By Lemma 3.11, there exist MCM $R$-modules $X$ and $Y$ such that $Y$ locally has finite injective dimension on the punctured spectrum, and $\operatorname{Ext}^{n+d}_{R}(M,N)\cong\operatorname{Ext}_{R}^{n}(X,Y)\cong\operatorname{Ext}_{R}^{n}(Y^{\dagger},X^{\dagger})$ for all $n\geqslant 1$, where the last isomorphism is by duality of dualizing module, see, e.g., [9, 3.14 and 3.15]. Since $Y$ is MCM and locally has finite injective dimension on the punctured spectrum, for every non-maximal prime ideal $\mathfrak{p}$ of $R$, the localization $Y_{\mathfrak{p}}$ is a finite direct sum of copies of $\omega_{\mathfrak{p}}$. Thus, $Y^{\dagger}$ is locally free on the punctured spectrum of $R$. Therefore, by [12, Lem. 3.6], there exists $h>0$ such that $0=\mathfrak{m}^{h}\operatorname{Ext}_{R}^{n}(Y^{\dagger},X^{\dagger})=\mathfrak{m}^{h}\operatorname{Ext}_{R}^{n+d}(M,N)$ for all $n\gg 0$. Hence, the proposition follows from Theorem 3.10. ∎
@@
In general, it may happen that $R$ is CM with isolated singularity, but $\widehat{R}$ is locally not even Gorenstein on the punctured spectrum, see [16, Main Theorem]. In the following remark, we note a class of local rings whose completions are locally Gorenstein on the punctured spectrum.
@@
    Suppose that $R$ is CM of dimension $1$. Then it is well known that $\widehat{R}$ is locally Gorenstein on the punctured spectrum if and only if $R$ admits a canonical ideal. Indeed, in view of [15, Prop. 2.7], $R$ admits a canonical ideal if and only if the total ring of fractions Q(R^)Q($\widehat${R}) is Gorenstein if and only if $\widehat{R}$ is locally Gorenstein at all minimal primes of $\widehat{R}$, where the last equivalence follows from [20, Lem. 10.25.4] and the fact that the direct product of finitely many rings is Gorenstein if and only if all the components are Gorenstein.
@@
First, assume that $\operatorname{G-dim}_{R}(X)<\infty$. Then, in view of 2.27, there exists an $R$-module $L$ of finite projective dimension and an injective homomorphism $h:X\to L$. By Lemma 2.19, $h(M)$ is a Burch submodule of $L$. Note that $\operatorname{Tor}^{R}_{\gg 0}(L,N)=0=\operatorname{Ext}^{\gg 0}_{R}(L,N)$, and $h(M)\cong M$. Therefore, we are done by Propositions 3.7 and 3.8.
@@
    The hypotheses $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(M,N)=0$ and/or $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(L,N)=0$ for some $h\geqslant 1$ in Propositions 3.1, 3.7 and Theorems 3.2, 3.15.(1) hold if $N$ locally has finite projective dimension on the punctured spectrum; see [12, Lem. 3.6.(2)].
@@
    The hypotheses $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,N)=0$ and/or $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,L)=0$ for some $h\geqslant 1$ in Propositions 3.3, 3.5, 3.9 and Theorems 3.6, 3.10, 3.15.(2) hold if $M$ locally has finite projective dimension on the punctured spectrum; see [12, Lem. 3.6.(1)].
@@
Although different in spirit from the rest of the results in this article, the following proposition demonstrates the usefulness of Lemma 2.19 by improving an injective dimension test result from [14, Cor. 5.4].
@@
(1) In view of 2.27, there exist an $R$-module $L$ of finite projective dimension and an injective homomorphism $h:X\to L$. By Lemma 2.19, $h(M)$ is a Burch submodule of $L$. Since $n>\operatorname{depth}(R)\geqslant\operatorname{pd}_{R}(L)$, it follows that $\operatorname{Ext}_{R}^{i}(L,N)=0$ for $i=n,n+1$. Since $h(M)\cong M$, by [14, Thm. 5.3], $\operatorname{id}_{R}(N)<\infty$.